TUNABLE INFRARED SPECTROMETER BASED ON DIFFERENCE-FREQUENCY GENERATION IN AgGaS_{2} FOR LINESHAPE STUDIES IN N_{2}O

\underline{A.VITCU}, R.WEHR, R.CIURYLO, J.R.DRUMMOND, A.D.MAY, Department of Physics, University of Toronto, ON, M5S 1A7, Canada.

A high resolution (2MHz) and high signal-to-noise ratio , cw, tunable, infrared spectrometer is presented. It is based on the difference-frequency generation technique and consists of combining two tunable ring dye lasers in a nonlinear crystal of AgGaS_{2}. The present output wavelength range is 7.8 to 9.1 microns. With total input power of 400 mW, the output power exceeds 100 nW at the peak of the emission curve. The spectrometer is currently used for line shape studies in the 03^{1}0 ← 01^{1}0 Q branch of N_{2}O, at 8.6 microns